OpenStudy (anonymous):

What is the range of the graph of y = -3(x - 4)2 + 1?

OpenStudy (paxpolaris):

anything squared has to be non-negative \[\implies \left( x-4 \right)^2 \ge0\]\[\implies -3\left( x-4 \right)^2 \le0\]\[\implies-3 \left( x-4 \right)^2+1 \le1\]

OpenStudy (paxpolaris):

\[\Large \therefore y\le1\]